OpenStudy (kfishburne23):

find slope intercept equation (-3,5) parallel to 2x-5y=6

OpenStudy (anonymous):

simplify the original eqt. -5y=-2x+6 y= 2/5x - 6/5 parallel means the 2nd equt. must have same slope = 2/5 y = mx+ b y = 2/5x+ b Now plug x and y in 5= (2/5)(/3) + b b= 19/5 or 3.8 so equt is y = 2/5x + 19/5

OpenStudy (anonymous):

whoops I messed up on the -3 part 5 = (2/5)(-3) + b b= 31/5 or 6.2 so answer is y= 2/5x + 31/5